OpenStudy (anonymous):

1.Find the volume of a pyramid which has a square base of 10 inches and a height of 24 inches.

OpenStudy (callisto):

Volume\[=\frac{1}{3} Area \times height = \frac{1}{3}10^2 \times 24 =?\]

OpenStudy (anonymous):

thats the answer?

OpenStudy (callisto):

You need to simplify it. Use your calculator please

OpenStudy (anonymous):

i got 800